What is the pH of a 1.50 mol/L solution of ammonium chloride (NH4Cl)

2. What is the pH of a 1.50 mol/L solution of ammonium chloride (NH4Cl)?

kb = 1.8 10^-5
Answer #1

Answer-

NH4Cl decomposes in water by following reaction -

NH4Cl \small \rightleftharpoons NH3 + HCl

According to the stoichiometry of the reaction-

1 mol of NH4Cl dissociates into 1 mol of NH3 and 1 mol of HCl

Let the volume of the solution = 1L

Then the initial moles of NH4Cl = 1.50

And let x moles of NH4Cl dissociates at equilibrium to forms x mol of NH3 and x mol of HCl

Now-

Time NH4Cl moles NH3 moles HCl moles
t = 0 1.50 mol 0 0
At equilibrium (1.50 - x) mol x mol x mol

Kb expression at equilibrium for this reaction is:

Kb = [NH3][HCl] / [NH3Cl]

The value of Kb is given as 1.8 x 10-5

\Rightarrow

1.8 x 10-5 = [NH3][HCl] / [NH3Cl]

1.8 x 10-5 = x.x / (1.50 - x)

1.8 x 10-5 = x2 / (1.50 - x)

Since NH4Cl is weak electrolyte here, the value of x will be very small

So we can say that, 1.50 - x \small \approx 1.50

Now

1.8 x 10-5 = x2 / 1.50

1.8 x 1.50 x 10-5 = x2

x2 = 1.8 x 1.50 x 10-5

x2 = 2.7 x 10-5

x2 = 27 x 10-6

x2 = \small \sqrt{27} x 10-3

x = 5.196 x 10-3

This means that 5.196 x 10-3 mol of HCl is formed at equilibrium

And we know that HCl is a strong electrolyte and completely dissociates to H+ and Cl- ions by following reaction-

HCl \small \rightarrow H+ + Cl-

Now according to the stoichiometry of the reaction-

1 mol of HCl forms \small \rightarrow mol of H+

5.196 x 10-3 of HCl forms \small \rightarrow 5.196 x 10-3 of H+

And the volume of the solution remains same (1L)

So the concentration of H+ ions = 5.196 x 10-3 mol/L

Now these free H+ ions will contribute for the pH of the solution, and the pH value is given by the relation-

pH = -log[H+]

pH = -log[ 5.196 x 10-3]

pH = 2.284

So the pH of this solution is 2.284